a base ball team consisting of 5 boys and 4 girls will be formed from a group of 6 boys and 7 girls. find how many different teams can be formed from the group?
Answer by greenestamps(13200)



You need to choose 5 of the 6 boys AND 4 of the 7 girls:

C(6,5)*C(7,4) = 6*35 = 210

ANSWER: 210

Note you chose "permutations" for the topic, but since nothing is said about the order in which they are picked, or about the positions each player plays, the problem is about combinations, not permutations.